### v2

# Decoherent histories with(out) objectivity in a (broken) apparatus

Abstract
We characterize monitored quantum dynamics in a solvable model exhibiting a phase transition between a measurement apparatus and a scrambler. We show that approximate decoherent histories emerge in both phases with respect to a coarse-grained extensive observable. However, the apparatus phase, where quantum Darwinism emerges, is distinguished by the non-ergodicity of the histories and their correlation with the measured qubit, which selects an ensemble of preferred pointer states. Our results demonstrate a clear distinction between two notion of classicality, decoherent histories and environment-induced decoherence.
How classical reality emerges from quantum mechanics is a fascinating yet ambivalent question: the term “classical” is semantically overloaded. The goal of this Letter is to distinguish the notion of classicality in two influential approaches: environment-induced decoherence, and decoherent histories, in a situation where their distinction is demonstrable and instructive.
The decoherence program [1, 2, 3, 4] is fundamentally motivated by a measurement problem, namely how to determine “what is being measured” when describing measurement as a unitary interaction between the system and the apparatus [5] . The key idea is that, the apparatus, and more generally, the environment) has many degrees of freedom, which allow to select a set of preferred “pointer states”, through the proliferation of classical records [6, 7]. In decoherence, and its refinement quantum Darwinism [8, 9, 10, 11, 12, 13, 14], classicality is defined as objectivity [15, 16, 17, 18], the redundant accessibility of records.
Meanwhile, the decoherent histories approach [19, 20, 21, 22, 23, 24] aims to describe classicality intrinsically, solely from the multi-time correlation of an isolated quantum system under unitary evolution. Classicality emerges if the quantum evolution can be well approximated by a stochastic sum over an ensemble of trajectories (the decoherent histories) in some configuration space, with vanishing interference between them.
Figure 1: A dynamically expanding tree model where every line is a qubit and every node is an isometry $v=v(\theta)$ (2). The root qubit $A$ is entangled with the qubit $S$ (1). The model has a transition at $\theta_{c}$ (7); it is an apparatus measuring $S$ if $\theta<\theta_{c}$, and a scrambler if $\theta>\theta_{c}$. In both phases, coarse-grained monitoring of the model (3),(13) yields decoherent histories (14) (Fig. 2). In the encoding phase, the histories are Ornstein-Uhlenbeck like (16) and uncorrelated with $S$. In the apparatus phase, the histories are non-ergodic (17) and selects an ensemble of pointer states of $S$ (see also Fig. 3).
Despite the intrinsic nature of classicality à la decoherent histories, many works have investigated decoherent histories in open quantum systems, motivated by the relation with the decoherence program [25, 26]. Some authors identify decoherent histories in the system (with the environment traced out) in the Markovian limit [27, 28], and make connection with quantum trajectories [29, 30, 31, 32]. Others treat the system and the environment as a whole, and derive decoherent histories (or a related branching structure) from objectivity [33, 34, 35]. This extensive literature fueled a folklore belief that decoherent histories require objectivity. This belief was challenged by recent works [36, 37, 38] showing evidence for (approximate) decoherent histories with respect to a coarse-grained and slowly varying observable, in an isolated macroscopic systems under chaotic evolution. Note that chaotic dynamics is known to scramble information, making them inaccessible for all practical purposes [39, 40, 41], which is the opposite of objectivity [42, 43]. Hence it seems that decoherent histories can exist independently of objectivity [44, 45, 46]. However the mechanisms appear entirely different, and the relation between the two notions of classicality awaits clarification.
In this Letter, we characterize decoherent histories in a solvable model [47, 48] where objectivity can be turned on or off: the model has a transition between a functioning apparatus and a scrambler (Fig. 1). We will show that decoherent histories emerge in both phases with respect to a coarse-grained observable, the coarse-graining being the essential common mechanism (Fig. 2). Nevertheless, we shall distinguish the phases by the (non)-ergodicity of the histories, and their correlation with the measured qubit, described in terms of a “pointer states ensemble” that we shall introduce (Fig. 3).
Model and objectivity threshold. Consider one of the simplest unitary models of measurement [5], where both the apparatus $A$ and the system $S$ are a qubit. Initially, they are disentangled and the apparatus is always in a “ready” state $|0\rangle_{A}$. The measurement unitary is such that $|i\rangle_{S}|0\rangle_{A}\mapsto|i\rangle_{S}|i\rangle_{A}$, for $i=0,1$, so that measuring the superposition state $(|0\rangle_{S}+|1\rangle_{S})/\sqrt{2}$ results in a maximally entangled pair:
$|\Psi\rangle=(|0\rangle_{S}|0\rangle_{A}+|1\rangle_{S}|1\rangle_{A})/\sqrt{2}\,.$ (1)
This is a situation where the aforementioned measurement basis problem is the most acute: it is impossible to know in which basis the measurement took place [6, 7].
In our apparatus model, $|\Psi\rangle$ is the initial state of a cascade process involving more and more qubits, bridging micro- and macroscopic realms [49, 50, 51]. Starting with the qubit $A$, at each time step, every existing qubit interacts with a new qubit (in a pure state), so that there are $N_{t}=2^{t}$ qubits after $t$ steps. During all this, $S$ remains intact. Such an inflationary dynamics is known as “concatenation” in the error correction code literature [52, 53, 54], and often represented diagrammatically by a tree, see Fig. 1 [55, 56]. Every node corresponds to an isometry that outputs two qubits from one, $v:\mathbb{C}^{2}\to\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, such that $v^{\dagger}v=I$. The evolution from time step $t$ to $t+1$ is given by $V_{t+1,t}=v^{\otimes 2^{t}}$. We also denote $V_{t,s}:=V_{t,t-1}V_{t-1,t-2}\dots V_{s+1,s}$ for $t\geq s$. We will focus on the following concrete family of models:
$v=\sum_{k=0,1}\left(e^{-iX\theta/4}|k\rangle\right)^{\otimes 2}\langle k|e^{-iX\theta/4}\,$, (2)
where $X$ is the Pauli-$x$ matrix, and $\theta\in(0,\pi/2)$ is the tuning parameter. Our model is equivalent to a quantum circuit with non-local control gates [57], similar to those realized by coherently transportable neutral atoms [58].
When $\theta=0$, our model realizes a repetition code, which clearly has objectivity: every single apparatus qubit has perfect classical correlation with the system qubit. Increasing $\theta$ perturbs the code, and, as we shall see, destroys the objectivity beyond some threshold $\theta_{c}$. To determine $\theta_{c}$, we consider the correlation between $S$ and an extensive observable of the apparatus, the total magnetization
$\mathcal{Q}_{t}=\sum_{j=1}^{N_{t}}Z_{j}.$ (3)
where $N_{t}=2^{t}$ and $Z_{j}$ acts on the $j$-th qubit at time $t$. At $\theta=0$, $\mathcal{Q}_{t}$ is also perfectly correlated with $Z_{S}$, the $Z$ operator on $S$. Exploiting the tree structure (see below and [57] for methods), we calculate the correlation between $O^{\prime}_{S}$ (for any $O^{\prime}$) and $\mathcal{Q}_{t}$ for any $\theta$. As $t\to\infty$, we find the following scaling behavior
$\mu_{t}:=\langle Z_{S}V^{\dagger}_{t,0}\mathcal{Q}_{t}V_{t,0}\rangle\sim N_{t}^{1-x_{1}}$ (4)
where $\left<[\dots]\right>=\left<\Psi|[\dots]|\Psi\right$>, and the exponent is
$x_{1}=-\log_{2}(\cos\theta).$ (5)
We also compute the scaling of the second moment of $\mathcal{Q}_{t}$:
$\eta^{2}_{t}:=$ $\langle V^{\dagger}_{t,0}\mathcal{Q}_{t}^{2}V_{t,0}\rangle=\langle V^{\dagger}_{t,0}\sum_{ij}Z_{i}Z_{j}V_{t,0}\rangle$
$\sim$ $\begin{cases}N_{t}&x_{1}>1/2\,,\\ N_{t}^{2(1-x_{1})}&x_{1}<1/2\,.\end{cases}$ (6)
The threshold corresponds to $x_{1}=1/2$ or
$\theta_{c}=\pi/4\,.$ (7)
When $\theta<\theta_{c}$, $x_{1}<1/2$, $\mu_{t}\sim\eta_{t}$, the correlation between $\mathcal{Q}_{t}$ and any operator on $S$ is comparable to its fluctuation amplitude. Moreover, (4) and (6) hold as well if $\mathcal{Q}_{t}$ (3) is replaced by the total magnetization of a fixed fraction $fN_{t}$ of apparatus qubits [57]. Thus, objectivity prevails: information about $S$ can be obtained from small fractions of the apparatus. We call $\theta<\theta_{c}$ the “apparatus phase”. When $\theta>\theta_{c},x_{1}>1/2$, the correlation is vanishing compared to the noise, $\mu_{t}\ll\eta_{t}$, and observing $\mathcal{Q}_{t}$ yields no information about the qubit. This is the “encoding phase”. Based on previous work on a very similar model [48], we expect that the Holevo quantity between an apparatus fraction and $S$ vanishes if $\theta>\theta_{c}$ as $t\to\infty$. In other words, information about $S$ is inaccessible to other observables as well, not just $\mathcal{Q}_{t}$.
We remark that our expanding tree model is a special case of the multi-scale entanglement renormalization Ansatz (MERA) tensor networks [59, 60, 61, 62], which are known to generate scale invariant states. In that context, the exponent $x_{1}$ above is the lightest scaling dimension. Interestingly, the threshold phenomena at $x_{1}=1/2$ appeared in various recently studied scale invariant “inference” problems [63, 64, 65].
Decoherent histories. We now turn to decoherent histories in our model. Our approach builds upon that of [37], which we first review. The standard decoherent histories formalism applied to our model would start with a family of projectors $\{K^{t}_{m}\}$ that sum to $\mathbf{1}$ for each time step $t\in[\tau,T]$. Then one defines, for $\vec{m}=(m_{\tau},\dots,m_{T})$, the state
$|\vec{m}\rangle:=K^{T}_{m_{T}}V_{T,T-1}\dots V_{\tau+1,\tau}K^{\tau}_{m_{\tau}}V_{\tau,0}|\Psi\rangle\,.$ (8)
Now, the norm of these states
$p_{\vec{m}}=\langle\vec{m}|\vec{m}\rangle$ (9)
are also the outcome probability distribution if one measures the apparatus at each time step projectively using $\{K_{m}^{t}\}$. The (exact) decoherent histories condition (DHC) is for these states to be orthogonal:
$\forall\,\vec{m}\neq\vec{m}^{\prime}\,,\,\left<\vec{m}|\vec{m}^{\prime}\right>=0\quad\text{(DHC)}$ (10)
A consequence of the DHC is that, if measurements are only performed at any time subset $t\in\mathbf{t}=\{t_{1}<t_{2}\dots<t_{k}\}\subset\{\tau,\dots,T\}$, the outcome distribution
$p^{\mathbf{t}}_{\vec{n}}=\|K_{n_{k}}^{t_{k}}V_{t_{k},t_{1}}\dots V_{t_{2},t_{1}}K_{n_{1}}^{t_{1}}V_{t_{1},0}|\Psi\rangle\|^{2}$ (11)
will be equal to the marginal of $p_{\vec{m}}$ to the time set $\mathbf{t}$:
$\text{DHC}\implies\Delta^{\mathbf{t}}_{\vec{n}}:=p^{\mathbf{t}}_{\vec{n}}-\sum_{\vec{m}|_{\mathbf{t}}=\vec{n}}p_{\vec{m}}=0\,$, (12)
where $\vec{m}|_{\mathbf{t}}=(m_{t_{1}},\dots,m_{t_{k}})$. Without DHC, $\Delta^{\mathbf{t}}_{\vec{n}}$ has no reason to vanish. In fact, outcome probability differences like $\Delta^{\mathbf{t}}_{\vec{n}}$ are essentially the only operational way to detect coherence. So it is reasonable to use the vanishing of (12) (for all $\textbf{t},\vec{n}$) as a probe of emergent approximate decoherent histories, as did the authors of Ref. [37].
In the sense of (12), classicality à la decoherent histories amounts to non-disturbance by third-party measurements (which is also part of macro-realism [66]). Namely, classicality emerges if the existence of measurements in $t\notin\mathbf{t}$ do not affect the outcome distribution of the ones in $t\in\mathbf{t}$. This is a reasonable way to define “classical” since only classical systems admit ideal measurements that reveal their state without changing it.
Given the above observation, we shall replace projective measurements by weak measurements described by smeared Kraus operators [67], which are more suitable for a coarse-grained observable. Concretely, we let
$K^{t}_{m}=\frac{1}{(2\pi\Gamma^{2})^{1/4}}e^{-(\mathcal{Q}_{t}/\eta_{t}-m)^{2}/(4\Gamma^{2})}\,,m\in\mathbb{R},$ (13)
where $\Gamma>0$ is $t$-independent. Namely, $K^{t}_{m}$ describes measuring the extensive observable (3) rescaled by its uncertainty (6), with $O(1)$ relative precision tuned by $\Gamma$. The idea of measuring the apparatus may evoke the Wigner friend’s paradox [68]. However, here, the “super-observer” is just a mathematical probe of the intrinsic dynamics of the apparatus (see also  [44, 45, 46, 69]).
Figure 2: Decoherent histories probe $\|\Delta^{\mathbf{t}}\|_{1}$ with $\mathbf{t}=\{L+\tau\}$, as function of $\tau$ (measurement starting time), for (a) $\theta=0.15\pi<\theta_{c}$ and (b) $\theta=0.3\pi>\theta_{c}$. Approximate decoherent histories emerge for coarse-grained measurements ($\Gamma=0.1$) and $\tau\gg 1$, but not for $\tau\sim 1$ or for fine-grained measurements (fg, $\Gamma\eta_{t}=0.1$). Insets (i, ii) show the pdf $p^{\mathbf{t}}$ (no third-party) and marginal distributions in non-decoherent scenarios ($T=11$). Straight lines show decay rate of $\|\Delta^{\mathbf{t}}\|_{1}$ (14), and (55) in [57]. (a) For $\theta<\theta_{c}$, $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{2\tau(x-1)}$ with no $L$ scaling. (b) For $\theta>\theta_{c}$, : $\|\Delta^{\mathbf{t}}\|_{1}\sim\Gamma^{-2}2^{-T}\max(1,2\cos\theta)^{2L}$. Inset (iii): $L$-dependence is shown for two $\theta>\theta_{c}$ ($\tau=6$).
It turns out that approximate decoherent histories emerge provided $\tau\gg 1$, that is, when the apparatus is macroscopic. We show in [57] that for any $\mathbf{t}$,
$\Delta_{\vec{n}}^{\mathbf{t}}=O((\Gamma\eta_{\tau})^{-2})\,,\tau\to\infty\,.$ (14)
Here we focus on the case $\mathbf{t}=\{T\}$, which has the most third-party measurements $t<T$ causing possible disturbance. Since the probe (12) involves a single variable, $n_{T}$, it is amenable to exact numerics. We evaluate the $L^{1}$ norm:
$\|\Delta^{\mathbf{t}}\|_{1}:=\int\left|\Delta^{\mathbf{t}}_{n_{T}}\right|dn_{T}\,$, (15)
whose vanishing indicates the emergence of decoherent histories. Some results are shown in Fig. 2. We observe that $\|\Delta^{\mathbf{t}}\|_{1}$ decays exponentially with respect to $\tau$ as predicted (14) for fixed $L=T-\tau$ in both phases. With respect to $L\gg 1$, there is essentially no dependence in the apparatus phase, while in the encoding phase, $\|\Delta^{\mathbf{t}}\|_{1}$ also decays exponentially. So, for small $\tau$, $\|\Delta^{\mathbf{t}}\|_{1}$ vanishes at large $T$ in the encoding phase and not in the apparatus phase. These results are explained as follows. Third-party measurements at the early, microscopic, stage will disturb the apparatus dynamics in both phases. In the apparatus phase, the disturbance remains accessible at late time. In the encoding phase, the disturbance is only visible at early time and becomes scrambled later.
The emergence of decoherent histories crucially relies on coarse-graining, that is, measuring an extensive quantity with $\Gamma=O(1)$ relative precision, see (13). Had we let $\Gamma\sim O(1/\eta_{t})$, $\|\Delta^{\mathbf{t}}\|_{1}\not\to 0$ in any limit (see blue crosses in Fig. 2). Such a measurement with $O(1)$ absolute precision would disturb the dynamics even in a macroscopic system, revealing its underlying quantum nature. Ref. [70] also pointed out the importance of measurement imprecision for classicality in macroscopic systems, in terms of the Leggett-Garg inequality [66]. Interestingly, in our model, this inequality can be violated at arbitrarily late time in both phases, by operators of type $Q_{t}=(aX_{j}+bY_{j}+cZ_{j})^{\otimes 2^{t}},Q_{t}^{2}=\mathbf{1}$ [57]. Again, such an observable reveals the many-body coherence of the model inaccessible to coarse-grained ones.
Histories statistics. Although decoherent histories emerge at $t\geq\tau\gg 1$ independently of objectivity, they have distinct statistics $p_{\vec{m}}$ (9) in the two phases. In the encoding phase ($x_{1}>1/2$), the time sequence $(m_{t})$ is a Gaussian process with finite temporal correlation:
$\mathbb{E}[m_{t}]=0,\mathbb{E}[m_{t}m_{t^{\prime}}]\sim e^{-\kappa|t-t^{\prime}|}\,,\,\kappa=(x_{1}-1/2)\ln 2$ (16)
for $|t-t^{\prime}|\gg 1$. The correlation time $\kappa^{-1}$ is finite (it only diverges near the threshold $x_{1}=1/2$). To understand this intuitively, recall that the extensive observable $\mathcal{Q}_{t\gg 1}$ reveals no initial information (stored in $S$); since the expanding dynamics is self-similar, information about the system at any time $t$ becomes inaccessible at $t^{\prime}-t\gg 1$ as well.
Meanwhile, the apparatus phase ($x_{1}<1/2$) is characterized by histories with the following probabilistic law:
$m_{t}\stackrel{{\scriptstyle\text{in law}}}{{=}}\mathcal{M}+\Gamma\,\xi_{t}$ (17)
where $\xi_{t}$ are i.i.d standard Gaussian and $\mathcal{M}$ is a random variable with the same distribution as $\mathcal{Q}_{t}/\eta_{t}$ in the long time limit (this distribution is non-Gaussian, see Fig. 2):
$\forall f\,,\,\overline{f(\mathcal{M})}=\lim_{t\to\infty}\langle V_{t,0}^{\dagger}f(\mathcal{Q}_{t}/\eta_{t})V_{t,0}\rangle.$ (18)
Namely, in every experimental run, the outcome $m_{t}$ becomes independent of $t$ up to an $O(1/\Gamma)$ noise, yet this constant value is random and varies at each run. Such behavior is akin to that of an apparatus pointer that freezes around a random value. The histories are non-ergodic: the statistical average over many runs differs from the time average over a single run. The correlation time is infinite: for any $t,t^{\prime}$, $m_{t}$ and $m_{t^{\prime}}$ are correlated. This is because $m_{t,t^{\prime}}$ is correlated with $\mathcal{Q}_{t,t^{\prime}}$ (respectively), and, as we have seen above, they are both correlated with the qubit $S$.
Figure 3: Distribution of pointer states and marginal outcome distribution $p_{m}$ for $\theta=0.05\pi<\theta_{c}$ (a,d), $\theta=0.15\pi<\theta_{c}$ (b,e) and $\pi=0.3\pi>\theta_{c}$ (c,f). The color code relates the outcome to the condition state of $S$ inside the Bloch sphere. The points are distributed uniformly with respect to the probability. ($t=20$ and $\Gamma=0.05$.)
Pointer states ensemble. We come back to the question: what information on the measured qubit $S$ is inferred from an apparatus history? This can be described by the ensemble of the conditional state $\rho_{\vec{m}}$ of $S$ after a coarse-grained monitoring [using (13)] of the apparatus, weighted by Born’s rule $p_{m}$ [48, 71, 72, 73], which we call the pointer states ensemble. This ensemble is different from and “softer” than the usual definitions of pointer states [6, 28, 74]. In general, it is made of mixed states that do not form a basis. Yet, they constitute a “decomposition of unity” like a basis: $\sum_{\vec{m}}p_{\vec{m}}\rho_{\vec{m}}=I$ [48]. Moreover, it captures an imperfect form of apparatus super-selection, as we now illustrate.
One may argue [57] that a full history of $t\in[\tau,T],\tau\gg 1$ infers no more information than a single-time measurement using (13) at $t\gg 1$, so we focus on the latter. Eq. (1) implies that the conditional density matrix of $S$ is
$\rho_{m}=(V_{t,0}^{\dagger}(K^{t}_{m})^{\dagger}K^{t}_{m}V_{t,0})^{\mathsf{T}}/(2p_{m})$ (19)
where $\mathsf{T}$ is the transpose. This can be efficiently computed, and has a well-defined $\Gamma\to 0,t\to\infty$ limit. Fig. 3 shows the ensemble of $\rho_{m}$ inside the Bloch sphere. The ensemble concentrates more and more on $(I\pm Z)/2$ as $\theta\to 0$. This is the limit of a perfect apparatus that super-selects the pointer state basis $\{|0\rangle,|1\rangle\}$. For $0<\theta<\theta_{c}$, we have an imperfect apparatus. There is a nonzero probability that the apparatus’ freezes at $m\sim 0$, while the qubit does not “collapse” towards either $|0\rangle$ or $|1\rangle$, but remains a superposition, $\rho_{m}\approx I+X$. In general, $\rho_{m}$ is mixed, meaning that the “collapse” is incomplete. Nevertheless, there is a temporal consensus [75] by virtue of (17): super-observers at different $t$ will agree with each other on what the apparatus did. In the encoding phase, $\rho_{m}\to I/2$: there is no meaningful pointer state.
Conclusion. We distinguished two notions of classicality. Defined via decoherent histories, its emergence in a macroscopic object relies only on coarse-graining and is likely generic. Yet, classicality à la quantum Darwinism is about the particular macro-microscopic correlation as realized during a quantum measurement. An apparatus is a macroscopic object, but not every macroscopic object is an apparatus. This statement may sound obvious, yet is nontrivial to demonstrate: we only did it in a somehow unusual “inflationary” model. Indeed, it would be desirable to confront objectivity and decoherent histories in conventional lattice models beyond small sizes. There, stabilizing objectivity is nontrivial, and plausibly requires some form of symmetry breaking [42, 76, 43, 51]. Meanwhile, implication of our study for the issue of classicality in cosmology [77] awaits exploration.
Acknowledgements.
We thank Serge Florens and Vincent Vennin for useful discussion and collaboration on related topics.